OpenStudy (anonymous):

log2(n) is: (a) ω( log10(n) ) (b) Θ( log10(n) ) (c) o( log10(n) )

OpenStudy (anonymous):

N+1= 2^-2 =1/4 so N=-3/4 is that log of (N+1) to the base 2? if yes, then 2^(-2)=N+1 using a scientific calculator (1/4)=N+1 N=(1/4)-1 N=-(3/4)